# Review of nonflow estimation methods and uncertainties in relativistic heavy-ion collisions

### v2

@@ 1 Introduction
Collective anisotropic flow is a hallmark of heavy-ion (nucleus-nucleus) collisions. It refers to azimuthal angular correlations among particles over the entire event – every particle is correlated with every other particle. One particular example is hydrodynamic flow, where the initial geometry anisotropy in finite impact parameter collisions is converted by interactions into momentum space anisotropy of particles [1]. All particles are thus correlated in azimuth to the ellipse-shaped collision geometry, e.g. with respect to the impact parameter direction. The anisotropy is nonzero even in head-on (zero impact parameter) collisions because of position fluctuations of nucleons inside the colliding nuclei giving rise to finite eccentricities [2, 3], which defines the so-called participant plane azimuth. Because of the same reason, nonzero anisotropy can also emerge in small-system collisions, such as proton-proton (pp), proton-nucleus (pA), deuteron-nucleus (dA), and helium-nucleus collisions [4, 5]. In fact, for any impact parameter, it is still the participant plane direction that is most relevant for azimuthal anisotropic flow, whose departure from that with respect to the impact parameter direction constitutes flow fluctuations, stemming out of fluctuations of the participant plane direction about the impact parameter direction [6].
The interactions in the system created in relativistic heavy-ion collisions at RHIC (Relativistic Heavy-Ion Collider) and the LHC (Large Hadron Collider), presumably the quark-gluon plasma (QGP), are governed by quantum chromodynamics (QCD) [7, 8, 9]. How exactly those interactions convert the initial-state geometry anisotropy into the final-state momentum anisotropy is not well settled. It is generally believed that ultra-strong interactions are required to produce the observed large anisotropy (or flow) in heavy-ion collisions, and the QGP created in those collisions is a nearly perfect fluid [10] and can be well described by hydrodynamics with little viscosity [11]. In peripheral heavy-ion collisions and small-system collisions, the interactions may not be intense enough where hydrodynamics could be applicable and the escape mechanism may be at work [12, 13, 14, 15]. It is important to keep in mind that the word “flow”, while suggestive, does not necessarily mean hydrodynamic flow.
@@
Here $n=1,2,3,...$; $v_{1}$ is called directed flow, $v_{2}$ elliptic flow, and $v_{3}$ triangular flow. Because $\psi_{n}$ is reconstructed from particles, Eq. (2) is equivalent to two-particle correlations. Namely, it can be obtained from two-particle cumulant, or the Fourier coefficient $V_{n}$ of the two-particle correlation function,
@@
where $v_{2}\{2\}$ denotes two-particle cumulant flow.
@@
If measured by the two-particle cumulant method of Eqs. (3,4), the $V_{n}$ unnecessarily must be related to the collision geometry. For example, the gluon field from each incoming nucleus can be correlated and this correlation may result in a final-state anisotropy $V_{n}$ [20, 21]. Such an initial-state correlation permeates over an entire collision event, so it is flow, but not part of hydrodynamic flow.
@@
There is, however, an important contribution to $V_{n}$ (or more generally, any variables) measured via correlations, and that is nonflow: two- and few-particle genuine correlations that have nothing to do with the collision geometry or event-wise global correlations [22, 23, 6]. Nonflow includes all other contributions to the measured $V_{n}$ except the collective flow. Examples of nonflow correlations are those between daughter particles from a resonance decay, particles from a jet shower originated from an energetic parton (quark or gluon) [24, 25], hadrons from string fragmentation [26], Hanbury-Brown Twiss (HBT) interferometry [27], and global momentum conservation [28]. Figure 1 illustrates those correlations in a two-particle angular correlation plot in $(\Delta\eta,\Delta\phi)$ where $\Delta\eta=\eta_{1}-\eta_{2}$ is the two-particle pseudorapidity difference. Most of the nonflow correlations (HBT, resonance decays, intra-jet correlations, string fragmentation) are small-angle correlations contributing to the near-side ($\Delta\phi\sim 0$) peak at $\Delta\eta\sim 0$. HBT is relevant at very small angle differences and generally considered minor. There is a long-range contribution from dijet correlations contributing to the away-side ridge at $\Delta\phi\sim\pi$ that has little $\Delta\eta$ dependence (because of the stochastic sampling of kinematics of the underlying parton-parton scattering producing the dijet). It is noteworthy that the nonflow contributions from jets are those intra- and inter-jet hadron correlations, whereas the azimuthal anisotropy of jet-axis orientations, a result of path-length dependent partonic energy loss or jet quenching [29, 30], is related to the collision geometry and is thus part of flow. Global momentum conservation contributes to the away-side ridge, but only to the first harmonic [28].
@@
Figure 1: Illustration of nonflow correlations in two-particle $(\Delta\eta,\Delta\phi)$ angular distribution. The near-side peak at $\Delta\phi\sim 0$ is mainly short ranged ($\Delta\eta\sim 0$). There could be various near-side short-range correlations: HBT is very short ranged (sharp peak), resonance decay and intra-jet correlations are short ranged with a $\Delta\eta$ width on the order of one unit. The away-side ridge at $\Delta\phi\sim\pi$ is long ranged and comes mainly from dijet correlations and global momentum conservation. Underlying these nonflow peaks are the majority pair distribution (zero-suppressed), modulated by anisotropic flows of various harmonic orders in $\Delta\phi$, which could have weak dependencies on $\eta$ (on single particle level) and $\Delta\eta$ (on two-particle level, referred to as flow decorrelation).
@@
All those nonflow correlations contribute to the $V_{n}$ of Eq. (4), so the factorization in Eq. (5) is no longer valid. A keen interest in relativistic heavy-ion collisions is to measure collective anisotropic flows arising from final-state interactions to probe the properties of the QGP by, for example, comparing to hydrodynamic calculations. Thus, the goal is to measure the collective anisotropic flows from global event-wise correlations related to the collision geometry. To this end, nonflow contamination in $V_{n}$ must first be subtracted with faithful systematic uncertainties [31]. Many experimental methods have been devised to estimate and/or subtract nonflow contributions. This note aims to give a pedagogical review of those methods in a single place, hopefully useful for future researches. The possible contributions from geometry-unrelated initial-state gluon correlations [20, 21] need also to be considered. This is however outside the scope of this note.
@@ 2.1 $\Delta\eta$-gap Methods
Nonflow correlations are primarily short ranged in $\Delta\eta$. Those short-range nonflow correlations can be suppressed by applying a $\Delta\eta$ gap between particle pairs used in the correlation analysis [16]. The $\Delta\eta$ gap cannot suppress those long-range nonflow correlations.
@@ 2.1.1 Simple $\Delta\eta$-gap method
The simplest way to suppress nonflow is to impose a $\Delta\eta$ gap between the two hadrons in two-particle cumulant measurement of $V_{n}$ [16]. This method is easy and straightforward. The shortcomings are obvious:
@@
    the away-side dijet correlations cannot be eliminated and how much they contribute to $V_{n}$ is unknown a priori;
@@
    $\eta$-dependent $v_{n}$ [32, 33] and longitudinal decorrelations [34, 35, 36] would yield different measurements with different $\Delta\eta$ gaps.
The last is not related to nonflow, but flow and flow fluctuations. However, since they cannot be distinguished by a $\Delta\eta$-gap analysis, the effects of flow and nonflow are mixed. All those shortcomings make the uncertainty estimation of a particular $\Delta\eta$-gap result difficult.
@@ 2.1.2 Two-subevent method
Instead of applying $\Delta\eta$-gap between the two particles, one may use the two-subevent method where one particle is taken from one subevent in a given $\eta$ region and the other from another subevent in a different $\eta$ region, and the two subevents are separated in $\eta$ with a certain $\Delta\eta$ gap [16]. This has the advantage that the cumulant calculation can be applied which involves only single-particle loops [37, 38], thus saving tremendous computing time compared to double loops of particle pairs. Obviously, all the shortcomings of the simple $\Delta\eta$-gap method described in Sect. 2.1.1 are present in the two-subevent method.
@@ 2.1.3 Three- and four-subevent methods
The away-side jet-correlations are long ranged in $\Delta\eta$, and thus the simple $\Delta\eta$-gap method and the two-subevent method cannot suppress the away-side nonflow correlations. One can use three- and four-subevent method to calculate multi-particle cumulants to suppress the away-side nonflow correlations from dijets by separating those subevents well in $\eta$ [39, 40, 41]. Figure 2 illustrates the idea behind the three-subevent method of four-particle cumulant, where dijets appear in at most two subevents and the third subevent is free of nonflow dijet correlations once the $\Delta\eta$-gaps are made wide enough. As a result, the correlations among the particles–two from one subevent and one from each of the other two subevents–are due only to collective flow. The four-subevent method for four-particle cumulant is similar, except that one particle is taken from each of the four subevents.
Figure 2: Sketch of the three-subevent method. Four-particle cumulant is formed by taking two particles from one subevent and one particle from each of the other two subevents. All nonflow correlations are suppressed including back-to-back dijet correlations, except global momentum conservation.
The multi-subevent method suppresses essentially all nonflow correlations (except that in $V_{1}$ from global momentum conservation [28]). This method also takes advantage of cumulant calculations [37, 38], greatly reducing the computing demand. The shortcomings are as same as those from the simple $\Delta\eta$-gap method described in Sect. 2.1.1, except (ii).
@@ 2.2 Low-multiplicity subtraction methods
The general idea behind this category of methods is that particle correlations in peripheral and low-multiplicity collisions are primarily due to nonflow, and those in central and high-multiplicity collisions come from collective flow with some contamination from nonflow correlations. The key questions are:
@@ 2.2.1 Inverse multiplicity scaling method
Resonance abundances are expected to scale approximately with the final-state multiplicity. If the scaling is exact and if resonance kinematic distributions, and thus the average correlations between decay daughters, do not vary with event multiplicity, then nonflow contribution from resonance decays to $V_{n}$ by Eq. (4) is inversely proportional to multiplicity ($N$). Take the number of particles to be $N_{\rm low}$ in low-multiplicity events and $N_{\rm high}$ in high-multiplicity events, and take the number of pairs to be $N^{2}$ (valid under Poisson statistics), then
$N^{2}_{\rm high}V_{2}^{\rm high}$ $=$ $N^{2}_{\rm high}V_{2}^{\rm sub}+N^{2}_{\rm low}V_{2}^{\rm low}\cdot\frac{N_{\rm high}}{N_{\rm low}}\,,$ (6)
$V_{2}^{\rm sub}$ $=$ $V_{2}^{\rm high}-\frac{N_{\rm low}}{N_{\rm high}}V_{2}^{\rm low}\,.$ (7)
@@
Here, it is assumed that the number of nonflow “sources” scales with multiplicity, and those nonflow particles, while keeping the nonflow correlations among themselves, have attained flow individually in high-multiplicity events. The notation $V_{2}^{\rm sub}$ denotes low-multiplicity subtracted $V_{2}$ with the ultimate goal to be the nonflow-subtracted anisotropic flow. This method is sometimes called the $c_{0}$ method where $c_{0}\equiv N$ [18, 19].
The $1/N$ scaling is a strong assumption for nonflow contribution from jet correlations:
@@
  • Jets are produced by hard processes whose abundance increases with $N$ more strongly than linearly [29, 30]. The $1/N$ scaling would be an underestimate of this part of nonflow. However, the majority nonflow contribution comes from relatively low transverse momentum ($p_{T}$) jets or minijets, whose production may be more closely proportional to $N$.
@@
  • Jets are modified by the nuclear medium created in relativistic heavy-ion collisions. Such modifications broaden/suppress the angular correlations between jet fragments [42, 43, 44], making nonflow weaker.
@@
  • On the other hand, jets lose energy via collisional and radiative partonic energy loss [45, 46], resulting in more lower $p_{T}$ particles and particle pairs, and thus stronger nonflow correlations.
@@
Quantitatively, these effects are $p_{T}$ dependent, and the interplay among them determines the final dependency of nonflow on collision centrality/multiplicity. Simulations by the hijing model (a jet production model without hydrodynamic flow) indicate only a modest excess of nonflow than the $1/N$ scaling by 10–20% from peripheral to central Au+Au collisions at $\sqrt{s_{\textsc{nn}}}=200$ GeV [31]. Nevertheless, this is one inferior part of the method.
@@
The $1/N$ scaling implies not only a scaling in the abundance of nonflow sources, but also that the physics of nonflow correlations does not change from low- to high-multiplicity collisions. In addition, the method explicitly assumes that the correlations in low-multiplicity collisions are all nonflow. The flow result in high-multiplicity events extracted from this method will thus inevitably depend on what low-multiplicity events are considered as the pure-nonflow baseline. The selection of very low-multiplicity or very peripheral events seems to be the natural choice, but they could be biased towards too soft underlying nucleon-nucleon scatterings or diffractive interactions. This is a selection bias [47, 48]. The nonflow in those events may not be a good reflection of nonflow in central collisions; in other words, the lowest multiplicity collisions are unnecessarily the best baseline for nonflow subtraction. This issue is more severe in analyzing anisotropies in small-system collisions, such as pp, pA, and dA collisions. Because of multiplicity selection biases [47, 48], the low- and high-multiplicity collisions of those small systems can be vastly different in physics, including nonflow correlations. High-multiplicity pp events are likely biased towards jet production, whereas low-multiplicity events are likely biased towards softer-than-average interactions.
To get a quantitative feeling, we examine Au+Au collisions at the top energy at RHIC. The measured $v_{2}$ in peripheral 70-80% and top 5% central Au+Au collisions are $v_{2}\{2\}^{\rm peri}\approx 6.9\%$ and $v_{2}\{2\}^{\rm cent}\approx 2.4\%$, respectively [49]. Assuming the former is all nonflow, the $1/N$ scaling of nonflow would result in a nonflow of $(v_{2}^{\rm peri})^{2}\frac{N_{\rm ch}^{\rm peri}}{N_{\rm ch}^{\rm cent}}\approx 1.7\times 10^{-4}$, i.e. 30% nonflow in $(v_{2}^{\rm cent})^{2}$ of the latter. Choosing 90-100% centrality as the low-multiplicity baseline would yield a smaller nonflow estimate [50]. However, it is possible that the 90-100% centrality, which is determined by event multiplicity, is biased towards too soft interactions, thus underestimating the nonflow to be used for central collisions. On the other hand, the higher 70-80% centrality data used above likely contain flow, and thus the 30% estimate is likely an overestimate of nonflow.
Multiplicity selection bias is likely insignificant in central heavy-ion collisions–there are many underlying nucleon-nucleon interactions. It is unlikely, up to certain high-multiplicity limit, to have all those interactions to produce jets, for example. Therefore, minimum-bias (MB) pp collisions would be the best baseline at our disposal for nonflow subtraction for central heavy-ion collisions within the scope of the $1/N$ scaling method. STAR has measured accumulative correlations ($NV_{2}$) as functions of $p_{T}$ in pp and Au+Au collisions (Fig. 1 in Ref. [51]). The ratio (pp/Au+Au) can be directly related to nonflow fraction in Au+Au under the assumption of $1/N$ scaling. The peripheral 80-100% Au+Au collisions are comparable to pp, suggesting that these peripheral collisions are dominated by nonflow. In the top 5% central collisions, the measured ratio is on average 15%, or 12% after correcting for the occupancy-dependent detector efficiency of 20% from pp to central Au+Au [52]. One may consider that the mean $\langle{p_{T}}\rangle$ in central Au+Au collisions is larger than in pp, and because $NV_{2}$ is a strongly increasing function of $p_{T}$, the $N$-weighted pp/Au+Au ratio (or the ratio of $NV_{2}$ at the corresponding $\langle{p_{T}}\rangle$’s) is smaller, approximately 8% [53]. However, since the multiplicity $N$ is already included in the $NV_{2}$ measure, this extra weighting may not be justified. On the other hand, one may argue that the pp/Au+Au ratio should be taken at higher $p_{T}$ for pp than for Au+Au because of jet energy loss in the latter, which would yield a larger ratio. These considerations suggest that nonflow estimations by the $1/N$ scaling have large uncertainties because of strong assumptions made in this method.
In Eq. (6), it is implicitly assumed that all particles in central events have attained the global collective flow by the $N_{\rm high}^{2}$ in front of $V_{2}$. This is reasonable for heavy-ion collisions but may not be justified in small-system collisions. It is possible that those nonflow particles, such as jets produced early, have exited the collision zone without participating in the final-state interactions that are responsible for the generation of flow. In such a case, the $V_{2}$ estimated by Eq. (7) would be an underestimate of flow in those high-multiplicity small-system collisions.
@@
To recap, the advantage of the $1/N$ scaling method is that it is easy and straightforward. The shortcomings are the inherent assumptions, namely,

### v4

@@ 1 Introduction
Collective anisotropic flow is a hallmark of heavy-ion (nucleus-nucleus) collisions [1]. It refers to azimuthal angular correlations among particles over the entire event – every particle is correlated with every other particle. One specific mechanism that gives rise to such global correlations (collective effects) is hydrodynamic flow, where the initial geometry anisotropy in finite impact parameter collisions is converted by interactions into momentum space anisotropy of particles [2, 3, 4]. All particles are thus correlated in azimuth to the ellipse-shaped collision geometry, e.g. with respect to the impact parameter direction. The anisotropy is nonzero even in head-on (zero impact parameter) collisions because of position fluctuations of nucleons inside the colliding nuclei giving rise to finite eccentricities [5, 6], which defines the so-called participant plane azimuth. Because of the same reason, nonzero anisotropy can also emerge in small-system collisions, such as proton-proton ($pp$), proton-nucleus (pA), deuteron-nucleus (dA), and helium-nucleus collisions [7, 8]. In fact, for any impact parameter, it is still the participant plane direction that is most relevant for azimuthal anisotropic flow, whose departure from that with respect to the impact parameter direction constitutes flow fluctuations, stemming out of fluctuations of the participant plane direction about the impact parameter direction [9, 10].
The interactions in the system created in relativistic heavy-ion collisions at RHIC (Relativistic Heavy-Ion Collider) and the LHC (Large Hadron Collider), presumably the quark-gluon plasma (QGP), are governed by quantum chromodynamics (QCD) [11, 12, 13, 14, 15, 16]. It is generally believed that ultra-strong interactions are required to produce the observed large anisotropy (or flow) in heavy-ion collisions, and the QGP created in those collisions is a nearly perfect fluid [17] and can be well described by hydrodynamics with little viscosity [18, 19, 20]. In peripheral heavy-ion collisions and small-system collisions, the interactions may not be intense enough where hydrodynamics could be applicable and the escape mechanism may be at work [21, 22, 23, 24]. Recently, it has been shown that collective effects can be generated with few interactions [23] and hydrodynamics is an attractor for the behaviour of many-body systems [25, 26, 27, 28]. It is important to keep in mind that the word “flow”, while suggestive, does not necessarily mean hydrodynamic flow.
@@
Here $n=1,2,3,...$; $v_{1}$ is called directed flow, $v_{2}$ elliptic flow, $v_{3}$ triangular flow, and so on. Because $\psi_{n}$ is reconstructed from particles, the $v_{n}$ values in Eq. (2) effectively measure two-particle correlation strengths. Namely, it can be obtained from two-particle cumulant, or the Fourier coefficient $V_{n}$ of the two-particle correlation function,
@@
where $v_{2}\{2\}$ denotes two-particle cumulant flow [33, 34, 35, 36].
@@
If measured by the two-particle cumulant method of Eqs. (3,4), the $V_{n}$ is not necessarily related to the collision geometry. For example, the gluon field from each incoming nucleus can be correlated and this correlation may result in a final-state anisotropy $V_{n}$ [37, 38]. Such an initial-state correlation permeates over an entire collision event, so it is flow, but not part of hydrodynamic flow.
@@
There is, however, an important contribution to $V_{n}$ (or more generally, any variables) measured via correlations, and that is nonflow: two- and few-particle genuine correlations that have nothing to do with the collision geometry or event-wise global correlations [39, 40, 41, 9, 42]. Nonflow includes all other contributions to the measured $V_{n}$ except the collective flow. Examples of nonflow correlations are those between daughter particles from a resonance decay, particles from a jet shower originated from an energetic parton (quark or gluon) [43, 44], hadrons from string fragmentation [45], Hanbury-Brown Twiss (HBT) interferometry [46], and global momentum conservation [47]. Figure 1 illustrates those correlations in a two-particle angular correlation plot in $(\Delta\eta,\Delta\phi)$ where $\Delta\eta=\eta_{1}-\eta_{2}$ is the two-particle pseudorapidity difference. Most of the nonflow correlations (HBT, resonance decays, intra-jet correlations, string fragmentation) are short-ranged (small-angle) correlations contributing to the near-side ($\Delta\phi\sim 0$) peak at $\Delta\eta\sim 0$. HBT is relevant at very small angle differences and its effect is generally minor because of the limited phase space at small angles. There is a long-range contribution from dijet correlations contributing to the away-side ridge at $\Delta\phi\sim\pi$ that has little $\Delta\eta$ dependence (because of the stochastic sampling of kinematics of the underlying parton-parton scattering producing the dijet) [48, 43]. It is noteworthy that the nonflow contributions from jets are those intra- and inter-jet hadron correlations, whereas the azimuthal anisotropy of jet-axis orientations, a result of path-length dependent partonic energy loss or jet quenching [49, 50], is related to the collision geometry and is thus part of flow. Global momentum conservation contributes to the away-side ridge, but only to the first harmonic [47].
@@
Figure 1: Illustration of nonflow correlations in two-particle $(\Delta\eta,\Delta\phi)$ angular distribution. The near-side peak at $\Delta\phi\sim 0$ is mainly short ranged ($\Delta\eta\sim 0$). There could be various near-side short-range correlations: HBT is very short ranged (sharp peak), resonance decay and intra-jet correlations are short ranged with a $\Delta\eta$ width on the order of one unit. The away-side ridge at $\Delta\phi\sim\pi$ is long ranged and comes mainly from dijet correlations and global momentum conservation. Underlying these nonflow peaks are the majority pair distribution (zero-suppressed), modulated by anisotropic flows of various harmonic orders in $\Delta\phi$, which could have dependencies on $\eta$ (on single particle level) and $\Delta\eta$ (on two-particle level, referred to as flow decorrelation).
@@
All those nonflow correlations contribute to the $V_{n}$ of Eq. (4), so the factorization in Eq. (5) is no longer valid [9, 51, 52, 53, 54]. A keen interest in relativistic heavy-ion collisions is to measure collective anisotropic flows arising from final-state interactions to probe the properties of the QGP by, for example, comparing to hydrodynamic calculations [55, 56, 20, 57]. Thus, the goal is to measure the collective anisotropic flows from global event-wise correlations related to the collision geometry. To this end, nonflow contamination in $V_{n}$ must first be subtracted with faithful systematic uncertainties [58]. Recently, azimuthal anisotropies similar to those in heavy-ion collisions have been observed in small-system p/d/3He+A and pp collisions [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 31, 32], prompting intense interest in possible collective phenomenon in small systems. There, nonflow correlations are dominating, if not the entirety of, the observed anisotropies. The challenge to remove nonflow correlations is particularly serious in order to identify any collective flow signals in small systems [69, 70, 71].
Many experimental methods have been devised to estimate and/or subtract nonflow contributions from azimuthal anisotropy measurements. This note aims to give a pedagogical review of those methods in a single place, hopefully useful for future researches. The possible contributions from geometry-unrelated initial-state gluon correlations [37, 38] need also to be considered. This is however outside the scope of this note.
It should be noted that nonflow is an issue mostly relevant to two-particle cumulant measurements as in Eqs (3,4,5). Nonflow is significantly reduced in multiparticle cumulants as clusters composed of several particles are sparse [72, 73, 74, 75, 76, 77]. Moreover, their contributions are diluted by multiplicity to higher powers [33, 34]. The flow fluctuation effects are, however, different in $v_{2}\{4\}$ and $v_{2}\{2\}$, negative in the former and positive in the latter [9, 78]. For physics pertinent to two-particle correlations, such as the chiral magnetic effect [79, 80], measurement of $v_{2}\{2\}$ is crucial and that of $v_{2}\{4\}$ is largely irrelevant. Furthermore, for rare probes, low-multiplicity events, or higher-order harmonics, $v_{n}\{4\}$ measurements are insurmountable statistically, so $v_{n}\{2\}$ measurements are essential. Because of all these reasons, nonflow studies in two-particle cumulant measurements are indispensable.
@@ 2 Nonflow estimation methods
There are three main categories of nonflow estimation methods: (1) applying pseudorapidity $\Delta\eta$ gaps between particle pair, (2) subtracting correlations from low-multiplicity events, and (3) performing data-driven fits to two-particle correlation functions.
@@ 2.1 $\Delta\eta$-gap Methods
Nonflow correlations are primarily short ranged in $\Delta\eta$. Those short-range nonflow correlations can be suppressed by applying a $\Delta\eta$ gap between particle pairs used in the correlation analysis. The $\Delta\eta$ gap method cannot suppress long-range nonflow correlations.
@@ 2.1.1 Simple $\Delta\eta$-gap method
The simplest way to suppress nonflow is to impose a $\Delta\eta$ gap between the two hadrons in two-particle cumulant measurement of $V_{n}$ [81, 82, 83, 84, 85, 86, 66]. This method is simple and straightforward to implement in data analysis. The shortcomings are obvious:
@@
    away-side jet correlations cannot be eliminated because dijets can be widely separated in $\eta$, and how much they contribute to $V_{n}$ is unknown a priori;
@@
    the $\eta$-dependencies [87, 73, 88, 89] and longitudinal decorrelations [90, 91, 92, 93] of flow and flow fluctuations would yield different measurements of $V_{n}$ with different $\Delta\eta$-gap sizes.
The last is not related to nonflow, but flow and flow fluctuations. However, since they cannot be distinguished by a $\Delta\eta$-gap analysis, the effects of flow and nonflow are mixed. These shortcomings make the uncertainty estimation of a particular $\Delta\eta$-gap result difficult.
@@ 2.1.2 Two-subevent method
Instead of applying $\Delta\eta$-gap between the two particles, one may use the two-subevent method where one particle is taken from one subevent in a given $\eta$ region and the other from another subevent in a different $\eta$ region, and the two subevents are separated in $\eta$ with a certain $\Delta\eta$ gap [94, 84, 95, 96, 97, 98]. This has the advantage that the cumulant calculation can be applied which involves only single-particle loops [36, 99], thus saving tremendous computing time compared to double loops of particle pairs. Obviously, all the shortcomings of the simple $\Delta\eta$-gap method described in Sect. 2.1.1 are present in the two-subevent method.
@@ 2.1.3 Three- and four-subevent methods
The away-side jet correlations are long ranged in $\Delta\eta$, and thus the simple $\Delta\eta$-gap method and the two-subevent method cannot suppress the away-side nonflow correlations. One can use three- and four-subevent method to calculate multi-particle cumulants to suppress the away-side nonflow correlations from dijets by separating those subevents well in $\eta$ [100, 97, 77]. Figure 2 illustrates the idea behind the three-subevent method of four-particle cumulant, where dijets appear in at most two subevents and the third subevent is free of nonflow dijet correlations once the $\Delta\eta$-gaps are made wide enough. As a result, the correlations among the particles–two from one subevent and one from each of the other two subevents–are due only to collective flow. The four-subevent method for four-particle cumulant is similar, except that one particle is taken from each of the four subevents.
Figure 2: Sketch of the three-subevent method. Four-particle cumulant is formed by taking two particles from one subevent and one particle from each of the other two subevents. All nonflow correlations are suppressed including back-to-back dijet correlations provided sufficient $\Delta\eta$ gaps, except global momentum conservation.
The multi-subevent method, with sufficient $\Delta\eta$ gaps, should suppress essentially all nonflow correlations (except those in $V_{1}$ from global momentum conservation [47]). This method also takes advantage of cumulant calculations [36, 99], greatly reducing the computing demand. The shortcomings are as same as those from the simple $\Delta\eta$-gap method described in Sect. 2.1.1, except (ii) provided sufficiently large $\Delta\eta$ gaps.
@@ 2.2 Low-multiplicity subtraction methods
The general idea behind this category of methods is that particle correlations in peripheral and low-multiplicity collisions are primarily due to nonflow, and those in central and high-multiplicity collisions come from collective flow with some contamination from nonflow correlations. Nonflow in $V_{n}$, being due to few-body correlations, is diluted by combinatorial particle pairs, and is therefore significantly reduced by high multiplicities. As a result, nonflow dominates in peripheral/low-multiplicity collisions and any flow may be neglected, and in central/high-multiplicity collisions flow may prevail. The degree of nonflow contamination at high multiplicities may then be assessed from particle correlation measurements in low-multiplicity events. The key questions are:
@@ 2.2.1 Inverse multiplicity scaling method
Resonance abundances are expected to scale approximately with the final-state multiplicity [101, 102]. If the scaling is exact and if resonance kinematic distributions, and thus the average correlations between decay daughters, do not change with event multiplicity, then nonflow contribution from resonance decays to $V_{n}$ by Eq. (4) is inversely proportional to multiplicity ($N$). Take the number of particles to be $N_{\rm low}$ in low-multiplicity events and $N_{\rm high}$ in high-multiplicity events, and take the number of pairs to be $N^{2}$ (valid under Poisson statistics), then
$N^{2}_{\rm high}V_{n}^{\rm high}$ $=$ $N^{2}_{\rm high}V_{n}^{\rm sub}+N^{2}_{\rm low}V_{n}^{\rm low}\cdot\frac{N_{\rm high}}{N_{\rm low}}\,,$ (6)
$V_{n}^{\rm sub}$ $=$ $V_{n}^{\rm high}-\frac{N_{\rm low}}{N_{\rm high}}V_{n}^{\rm low}\,.$ (7)
@@
Here, it is assumed that the number of nonflow “sources” scales with multiplicity, leading to the factor $N_{\rm high}/N_{\rm low}$ in Eq. (6). The notation $V_{n}^{\rm sub}$ denotes low-multiplicity subtracted $V_{n}$ with the ultimate goal to be the nonflow-subtracted anisotropic flow. This method is sometimes called the $c_{0}$ method where $c_{0}\equiv N$ [31, 32].
The $1/N$ scaling implies not only a scaling in the abundance of nonflow sources, but also that the physics of nonflow correlations does not change from low- to high-multiplicity collisions. Neither is good assumption for nonflow contribution from jet correlations:
@@
  • Jets are produced by hard processes whose abundance increases with $N$ more strongly than linearly in heavy-ion collisions [49, 50]. The $1/N$ scaling would be an underestimate of this part of nonflow. However, the majority nonflow contribution comes from relatively low transverse momentum ($p_{T}$) jets or minijets, whose production may be more closely proportional to $N$.
@@
  • Jets are modified by the nuclear medium created in relativistic heavy-ion collisions. Such modifications broaden/suppress the angular correlations between jet fragments [103, 104, 105, 106, 107, 44, 108], making nonflow weaker.
@@
  • On the other hand, jets lose energy via collisional and radiative partonic energy loss [48, 109, 110], resulting in more lower $p_{T}$ particles and particle pairs, and thus stronger nonflow correlations.
@@
Quantitatively, these effects are $p_{T}$ dependent, and the interplay among them determines the final dependency of nonflow on collision centrality/multiplicity. Simulations by the hijing model–a jet production model without hydrodynamic flow so its correlations are entirely of nonflow nature [^1] The pathlength dependent energy loss results in an anisotropy of jet fragments relative to the overall collision geometry, thus flow. This flow is encoded in the correlations between jet fragments and underlying particles in hijing but is negligible compared to the nonflow caused by intra- and inter-jet correlations [111].–indicate only a modest excess (10–20%) of correlation strengths in central Au+Au collisions at $\sqrt{s_{\textsc{nn}}}=200$ GeV than the $1/N$ scaling from peripheral collisions would entail [58]. Nevertheless, this is one inferior part of the method.
@@
Besides the strong assumptions, the $1/N$ scaling method also explicitly assumes that the correlations in low-multiplicity collisions are all nonflow. The flow result in high-multiplicity events extracted from this method will thus inevitably depend on what low-multiplicity events are considered as the pure-nonflow baseline. The selection of very low-multiplicity or very peripheral events seems to be the natural choice, but they could be biased towards too soft underlying nucleon-nucleon scatterings or diffractive interactions. This is a selection bias and has indeed been observed in small systems at RHIC [112, 113] and in peripheral Pb+Pb collisions at the LHC [114]. Such effects at the LHC start to become important at 80% centrality mark towards more peripheral collisions [114]. The nonflow in those events may therefore not be a good reflection of nonflow in central collisions; in other words, the lowest multiplicity collisions are unnecessarily the best baseline for nonflow subtraction.
This issue is more severe in analyzing anisotropies in small-system collisions, such as $pp$, $p$A, and $d$A collisions. Because of multiplicity selection biases [112, 113], the low- and high-multiplicity collisions of those small systems can be vastly different in physics, including nonflow correlations. High-multiplicity $pp$ events are likely biased towards jet production, whereas low-multiplicity events are likely biased towards softer-than-average interactions. As a result, the $1/N$ scaling may completely fail in gauging nonflow contamination in small systems.
Multiplicity selection bias is likely insignificant in central heavy-ion collisions–there are many underlying nucleon-nucleon interactions. For example, it is highly unlikely to have all those interactions to produce jets. [^2] This is of course valid only up to a certain high-multiplicity limit. When one demands high multiplicity with extremely small cross section, multiplicity selection biases would also manifest in central heavy-ion collisions. Therefore, minimum-bias $pp$ collisions would be the best baseline at our disposal for nonflow subtraction for central heavy-ion collisions within the scope of the $1/N$ scaling method.
In Eq. (6), it is implicitly assumed, by the $N_{\rm high}^{2}$ in front of $V_{n}^{\rm sub}$, that those nonflow particles have attained the global collective flow with all other particles in central events, while still retaining their genuine few-body correlations. This can happen in two scenarios: the particles have attained collective flow from system expansion but still retaining their correlations from early times, or the correlations are generated late in time after the collective flow has built up. These are reasonable for heavy-ion collisions but may not be justified in small-system collisions. It is possible that those nonflow particles, such as jets produced early, have exited the collision zone without participating in the final-state interactions that are responsible for the generation of flow. In such a case, the $V_{n}$ estimated by Eq. (7) would be an underestimate of flow of the underlying event in those high-multiplicity small-system collisions, and that from the template fit method (to be discussed in Sect. 2.2.3) may be a better reflection of the underlying flow. Another good example is resonance decays, where any flow attained by the parent particle is inherited in the decay daughters while the nonflow correlations between the daughter particles (dependent of the decay kinematics) still contribute to the final-state two-particle anisotropy. This applies to both heavy-ion and small-system collisions, and therefore it makes more sense to count all particles, including those decay products, in the baseline to evaluate azimuthal anisotropies. How accurate this part of nonflow is estimated will of course depend on how the abundances and kinematic distributions of resonances vary with the event multiplicity.
@@
To recap, the advantage of the $1/N$ scaling method is that it is simple to implement. The shortcomings are the inherent assumptions, namely,